# UCIL: An Unsupervised Class Incremental Learning Approach for Sound Event Detection

### v1

@@ Abstract
This work explores class-incremental learning (CIL) for sound event detection (SED), advancing adaptability towards real-world scenarios. CIL’s success in domains like computer vision inspired our SED-tailored method, addressing the unique challenges of diverse and complex audio environments. Our approach employs an independent unsupervised learning framework with a distillation loss function to integrate new sound classes while preserving the SED model consistency across incremental tasks. We further enhance this framework with a sample selection strategy for unlabeled data and a balanced exemplar update mechanism, ensuring varied and illustrative sound representations. Evaluating various continual learning methods on the DCASE 2023 Task 4 dataset, we find that our research offers insights into each method’s applicability for real-world SED systems that can have newly added sound classes. The findings also delineate future directions of CIL in dynamic audio settings.
Index Terms—  continual learning, sound event detection
1 Introduction
Sound event detection (SED) [1, 2] focuses on the automatic identification and classification of various sound events within an audio stream. This technology underpins a wide array of applications, from enhancing urban security systems through the detection of anomalous noises to monitoring wildlife habitats by recognizing species-specific calls [3, 4, 5]. The essence of SED lies in its ability to parse continuous audio input, segment it into discrete events, and accurately label each frame with one or more predefined sound classes. Advances in deep learning have significantly propelled the capabilities of SED systems, enabling them to perform effectively in controlled environments. These systems generally operate under a closed set framework, where the model is trained on a fixed dataset comprising predetermined sound classes. However, there can be possibilities, where there emerges a need to deal with recognition of new sound events from the perspective of the end-user for real-world scenarios.
@@ 1 Introduction
In response to the evolving nature of sound events in practical applications, different strategies have been explored to incorporate new sound classes into existing SED models. The most common approach is fine-tuning [6, 7], where the model pretrained on the original set of classes is further trained using a smaller dataset containing the new classes. While fine-tuning can facilitate the integration of new knowledge [8, 9, 10], it often leads to catastrophic forgetting, that affects performance of the model for the previously learned classes as it accommodates new information. This limitation leads to a challenge where the model struggles to learn new classes without losing accuracy on the ones it already knows. Essentially, as the model tries to absorb new information, its performance in previously learned classes degrades significantly. The challenge is not just about introducing new classes; it’s about enabling the model to expand its abilities over time while still holding onto the knowledge it initially acquired.
The above discussed backdrop sets the stage for continual learning [11, 12] that offers a novel paradigm designed to overcome limitations of a model learned on a closed set of classes to continuously acquire new knowledge while preserving existing knowledge. This approach counters catastrophic forgetting, thereby ensuring that models adapt effectively over time. Class-incremental learning (CIL) [13] is distinguished by its capacity to integrate new classes into a model’s learning architecture sequentially. Unlike task-incremental learning, which employs separate classifiers for each task, CIL utilizes an expanding classifier, facilitating the model’s growth as new data becomes available without the need for a full retraining.
CIL is pivotal in retaining previously learned knowledge, proving essential for applications requiring ongoing model evolution. It has been effective in areas such as computer vision (CV) [14, 15] and natural language processing[16], significantly contributing their advancements. Its extension into the audio [17] and speech domain [18], covering tasks like acoustic scene classification (ASC) [19, 20] and keyword spotting[21], has proven the method’s adaptability and utility in managing dynamic datasets. However, its application in SED is relatively recent. While Koh et al. [22] have made strides in adapting models for SED through transfer learning, it leaves the scope for incremental learning to develop real-world SED systems.
@@
This work introduces a novel CIL method for SED that aims to incorporate new sound categories and maintain accurate detection of sound events. Drawing inspiration from techniques used in ASC [19, 20] and CV [23, 24], we implemented an unsupervised class incremental learning (UCIL) framework to efficiently train the classifier on different tasks. We integrate a distillation loss function that helps in retaining the classifier’s knowledge of previously learned sound events while learning new ones, minimizing feature differences between model iterations. Our strategy also includes a sample selection process to leverage unlabeled data for improving the consistency of SED model.
@@
Further, we propose a technique to preserve previous examples by considering the duration and frequency of events within each sound category ensuring a balance. The proposed UCIL is evaluated on the DCASE 2023 Task 4A dataset with comparison to some CIL methods. This work aims to contribute in the ongoing development of CIL for audio processing considering SED as an application, indicating future research directions for real-world scenarios amidst dynamic audio settings.
2 Class Incremental Learning
2.1 Class Incremental Learning for SED
@@
We introduce a CIL approach for SED, where the model learns to recognize a variety of sounds expanded through a series of tasks $\tau_{0},\tau_{1},\ldots,\tau_{N}$ each representing a different set of sounds to be identified. This learning journey unfolds over $1+N$ time steps, starting with an initial task $\tau_{0}$ and moving through $N$ subsequent tasks, each introducing new sounds for detection. Every task $\tau_{i}$ has its own dataset $X_{i}$, consisting of sound features $x_{j}^{i}$ and their corresponding labels $y_{j}^{i}$ that are represented as multi-hot vectors. The total number of sound classes $|C|$ evolves to include both existing classes learned in previous tasks ($|C_{\text{ext}}|$) and the new classes introduced in the current task ($|C_{\text{new}}|$).
Our model, designated as $P_{i}$ for each task $\tau_{i}$, comprises two main components: a feature extractor $F_{i}^{\theta}$ and a classifier $H_{i}^{\phi}$. The feature extractor analyzes the sound data to identify distinguishing characteristics, while the classifier uses these features to predict the types of sounds present.
2.2 Proposed UCIL Method
@@
2.2.1 Independent Learning to Update Model by New Data
In our framework for SED, we start by training a learning model on a fundamental task, which we refer to as $\tau_{0}$. This initial task uses a dataset named $X_{0}$. To measure how well our model is learning, we employ binary cross-entropy loss ($L_{BCE}$) that computes the difference between the model predictions and the ground truth during both the initial learning stage and later stages. It operates by applying a mathematical function, the sigmoid function $\sigma$, to the model’s logits $o_{k}$ in class $k$ as:
$L_{BCE}=-\frac{1}{|C|}$ $\sum_{k=|C|}^{|C|}y_{k}^{i}\cdot\log(\sigma(o_{k}))$ (1)
@@ 2.2.1 Independent Learning to Update Model by New Data
After the initial task $\tau_{0}$, introducing the model to new sounds in subsequent tasks $\tau_{i}$, we create an updated version of our model denoted as $P_{i}$. We achieve this by adding the new classification heads of the linear layer, building on the previous model $\hat{P}_{i-1}$ to accommodate the new sound classes $|C_{\text{new}}|$ of the current task. In other words, $P_{i}$ is an expanded version of $\hat{P}_{i-1}$, equipped with these new classifiers. When we update the model, we focus only on enlarging its classification ability to include the new sounds. The predictions made by the model are now a mix of existing and new sound classes, which we refer to as $o_{ext}$ and $o_{new}$, respectively. The updated model, with its new feature analyzer $\hat{F}_{i}^{\theta}$ and classifier $\hat{H}_{i}^{\phi}$, is then trained on the latest data set, $\hat{X}_{i}$, to handle the current SED task. To prevent data imbalance issues of existing and new classes, we use independent learning (IndL) [20] method, which allows the model to learn the new sound predictions independently from the existing ones. Specifically, we only compute the $L_{BCE}$ by applying the sigmoid function $\sigma$ to only the new predictions. This way, the existing predictions are maintained separately, helping the model to remember and recognize both existing and new sounds effectively.
2.2.2 Temporal Knowledge Distillation from Existing to New
@@
We use two special distillation losses to maintain the learned knowledge of our model. The first one is output distillation loss $\mathcal{L}_{OD}$, which reduces differences in the model’s initial predictions for known sound classes, using mean squared error (MSE) loss. The second one is the feature distillation loss $\mathcal{L}_{FD}$, which minimizes the gap between the current and existing model’s feature extractor outputs by comparing their alignment with cosine similarity. The current model $P_{i}$’s feature is the output of the feature extractor $F_{i}^{\theta}$ given as $v=F_{i}^{\theta}(x)$. In simpler terms, these techniques help our updated model $P_{i}$ act like its existing version, $\hat{P}_{i-1}$, especially for sounds it already learned.
$\mathcal{L}_{FD}=1-\cos(\text{norm}(v),\text{norm}(\bar{v}))$ (2)
$\mathcal{L}_{OD}=MSE(\hat{P}_{i-1}(x)||o_{ext})$
$\mathcal{L}_{KD}=\mathcal{L}_{FD}+\lambda\mathcal{L}_{OD}$
@@
where norm and $\cos$ denote the L2-normalization and cosine similarity, respectively. In detail, we compute the scalar product with normalized vectors to get the cosine similarity. The weight $\lambda=\Omega\sqrt{\frac{|C|}{|C_{new}|}}$ is set adaptively, $\Omega$ is a fixed constant, and $\lambda$ increases as the number of incremental phases increases, to retain the knowledge of the increasing number of existing classes in the model. For the existing sounds, we measure the output predictions from $P_{i}$ and match them against $\hat{P}_{i-1}$.
For the new sound classes, we ensure that the model learns to extract consistent features with the previous model. The model’s training involves a mix of these methods, with a carefully tuned balance, to keep the model’s knowledge up-to-date without forgetting existing information.
@@
Figure 1: Block diagram of the proposed UCIL approach for task $\tau_{i}$. All training data for task $\tau_{i}$ include the new data $X_{i}$, the rehearsal data $x^{e}_{i}$, and the unlabeled data $x^{u}_{i}$. Bottom-right corner shows how different data are used to predict different outputs.
2.2.3 Unsupervised Learning with Sample Selection
The essence of model distillation above in CIL is to ensure the capability for the new version of a model is not too deviated from what it has learned before. It resembles to the scenario of trying to make sure that a student’s new knowledge does not overwrite what they already know. A common method to do this involves some assumptions about the new and previous knowledge never overlapping, but these assumptions often do not hold up in real-world scenarios that leads the model to forget the previous knowledge [25].
[…]
$d=|\hat{P}_{i-1}(x^{u}_{i})-P_{i}(x^{u}_{i})|$ (3)
@@ 2.2.3 Unsupervised Learning with Sample Selection
In practice, we take the first half of samples based on the discrepancy $d$ from our additional unlabeled data for each batch of training task $\tau_{i}$. By selecting these samples by discrepancy carefully, we assist the model to retain previously learned knowledge and become generalized from the past knowledge to new situations.
We adapt $\mathcal{L}_{OD}$ to use as $\mathcal{L}_{UOD}$, focusing on unlabeled data rather than the labeled data for $\mathcal{L}_{OD}$ as given in Section 2.2.2 following our goal to reduce the gaps between the new and existing models using this unlabeled data.
2.2.4 Balanced Memory Update Method
@@
To further maintain the previous knowledge, especially in applications with complex classes such as SED, we preserve some of the data from existing classes in advance as the rehearsal data. We mix this data $x^{e}_{i}$ with the incoming data of task $\tau_{i}$. The balanced sampling function is an algorithm tailored for preparing rehearsal data in the context of SED using strongly and weakly labeled data. It starts by computing the duration of sound events, considering the difference between their onset and offset times. Then, it aggregates these durations for each event, providing a measure of total exposure for each sound in the strongly labeled dataset. For weakly labeled data, the duration is 10 seconds for each event label.
@@
The method makes sure if there is a fair mix of sounds for training the model. It figures out how many examples of each sound are needed based on the seen classes. It then picks a balanced set of sound examples from both types of data to avoid any bias. The outcome is two rehearsal datasets, one from strongly labeled and the other from weakly labeled data. This way, all seen sounds get a fair chance to replay during training, leading to a more reliable and fair SED model.
@@
We adopt a multifaceted approach for independent unsupervised learning to handle the challenge of integrating new sound classes without forgetting previously learned information as illustrated in Figure 1. The total loss function of the updated model is given as:
$\mathcal{L}_{tot}=\mathcal{L}_{BCE}+\mathcal{L}_{KD}+\mathcal{L}_{UOD}$ (4)
@@
In summary, the model is initially trained on a foundational task and then updated through subsequent tasks with new classifier units to enhance its classification capabilities, employing independent learning to segregate the learning of new and existing sound classes. To preserve accumulated knowledge of the model, we use output and feature distillation losses, minimizing discrepancies between the model versions and ensuring consistency across all sound classes. Additionally, we incorporate unsupervised learning with strategic sample selection from unlabeled datasets to expose the model to a broader range of data, thus improving its generalization capabilities. Lastly, our balanced memory update method prepares rehearsal data from both strongly and weakly labeled datasets, ensuring a fair representation of all sound classes to mitigate data bias.
3 Experiment Setting
3.1 Dataset, Task Setting and Performance Metric
@@
In this study, we use the DCASE 2023 Task 4A dataset [26], a comprehensive collection of 10-second audio clips categorized into 10 sound events. It comprises of 1,578 real recordings with weak annotations, 14,412 real recordings without labels, and 10,000 synthetic recordings that are strongly labeled [27]. All clips were resampled to 16 kHz mono and segmented using a 2048-sample window and 256-sample hop length for spectrogram extraction and log-mel spectrogram generation.
@@
We devised two incremental task settings to assess our proposed method. Firstly, the two-task setting randomly divides the 10 target sound classes in the dataset into two groups, each containing 5 classes. This approach is designed to test how well our method can adapt from one group of sounds to another, a process known as source-target domain adaptation in SED. Secondly, the four-task setting is inspired by an existing multi-task learning strategy [28], that organizes the 10 target classes into 4 groups based on their acoustic characteristics: [vacuum cleaner, frying, blender], [electric shaver/toothbrush, running water], [speech, dog, cat], and [dishes, alarm bell ringing]. We shuffle the groups as the four tasks to investigate the effect of the number of tasks on the model’s ability to distinguish between different sounds, providing insights into the complexities of learning and recognizing diverse acoustic events.
@@
3.2 Implementation Details and Reference Baselines
In our experiments, we used a batch size of 48 and the Adam [31] optimizer with an initial 0.001 learning rate, gradually increasing over the first 50 of 200 total epochs following DCASE 2023 Task 4A setup. We adopted a mean teacher model strategy [32] to enhance training stability, applying an exponential moving average with a factor of 0.999. We consider a convolutional recurrent neural network (CRNN) [33] model that serves as the DCASE baseline with about 1.2M parameters for all the studies in this work. We implemented early stopping at 50 epochs to prevent overfitting, and set the weight $\Omega$ for the output distillation loss ($L_{OD}$) at 2, ensuring a balance between retaining existing knowledge and learning new information. We built following 5 reference baselines:
Table 1: Performance comparison of our proposed UCIL method with other CIL methods using the two-task settings. Rehearsal size indicates the maximum clip numbers of [Strongly Labeled, Weakly Labeled] as the rehearsal data in each task.

### v2

@@ Abstract
This work explores class-incremental learning (CIL) for sound event detection (SED), advancing adaptability towards real-world scenarios. CIL’s success in domains like computer vision inspired our SED-tailored method, addressing the unique challenges of diverse and complex audio environments. Our approach employs an independent unsupervised learning framework with a distillation loss function to integrate new sound classes while preserving the SED model consistency across incremental tasks. We further enhance this framework with a sample selection strategy for unlabeled data and a balanced exemplar update mechanism, ensuring varied and illustrative sound representations. Evaluating various continual learning methods on the DCASE 2023 Task 4 dataset, our research offers insights into each method’s applicability for real-world SED systems that can have newly added sound classes. The findings also delineate future directions of CIL in dynamic audio settings.
Index Terms:
continual learning, sound event detection
I Introduction
Sound event detection (SED) [1, 2] identifies and classifies various sounds within an audio stream. It supports applications like urban security and wildlife monitoring by detecting specific noises [3, 4, 5]. SED parses continuous audio, segments it, and labels each event with predefined sound classes. Deep learning has greatly improved SED, making it effective in controlled environments. These systems generally operate under a closed set framework with a fixed dataset comprising predetermined sound classes but real-world scenarios require incrementally recognizing new sound events [6].
@@ I Introduction
In response to the evolving nature of sound events in practical applications, different strategies have been explored to incorporate new sound classes into existing SED models. The most common approach is fine-tuning [7, 8], where a pretrained model is retrained using a smaller dataset of new classes. While fine-tuning helps integrate new knowledge [9, 10, 11], it can negatively impact the model’s performance in previously learned classes due to catastrophic forgetting. This challenge arises because the model struggles to learn new classes without losing accuracy on the ones it already knows. The practical issue is not just about introducing new classes; it’s about enabling the model to expand its abilities over time while still holding onto the knowledge it initially acquired.
The above context highlights the need for continual learning [12, 13], a paradigm that enables models to continuously acquire new knowledge while preserving existing information. This approach addresses catastrophic forgetting, ensuring models can adapt and evolve effectively over time. In continual learning, class-incremental learning (CIL) [14] integrates new classes into a model’s learning architecture sequentially. CIL is pivotal in retaining previously learned knowledge, proving essential for applications requiring ongoing model evolution. It has been effective in areas such as computer vision (CV) [15, 16] and natural language processing[17], significantly contributing to their advancements.
For CIL’s extension into the audio [18] and speech domain [19], covering tasks like audio classification (AC) [20, 21, 22] and keyword spotting[23], has proven the method’s adaptability and utility in managing dynamic datasets. Inspired with [24, 25], Mulimani et al. [20, 22] introduce one independent learning method with two distillation losses for the CIL of multi-label audio classification. However, their approach is for audio classification with clip-wise labels and does not compare with other continual learning methods. Although strongly labeled SED has been explored much in recent years under the DCASE challenge Task 4 [26], the CIL application in SED with strong labels to estimate the precise onset and offset is relatively recent. While Koh et al. [27] have made strides in adapting models for SED through transfer learning, it does not cover the scope for incremental learning to develop real-world strong labeled SED systems.
@@
This work introduces a novel CIL method for SED that aims to incorporate new sound categories and maintain accurate detection of sound events. Drawing inspiration from techniques used in [24, 20], we implemented an unsupervised class incremental learning (UCIL) framework to efficiently train the classifier on different tasks. We integrate a distillation loss function that helps in retaining the classifier’s knowledge of previously learned sound events while learning new ones, minimizing feature differences between model iterations. Because the strong label is limited, our strategy also includes a sample selection process to leverage unlabeled data for improving the consistency of the SED model.
@@
Further, we propose a technique to preserve previous examples for future rehearsal by considering the duration and frequency of events within each sound category ensuring a balance. The proposed UCIL is evaluated on the DCASE 2023 Task 4A dataset [26] with comparison to some CIL methods. To the best of our knowledge, this is the first work to apply rehearsal-based CIL techniques to the strongly labeled SED, thereby opening the door for future research directions in real-world scenarios for SED applications.
II Class Incremental Learning
II-A Class Incremental Learning for SED
@@
We introduce a CIL approach for SED, where the model learns to recognize a variety of sounds expanded through a series of tasks $t_{0},t_{1},\ldots,t_{T}$ each representing a different set of sounds to be identified. This learning journey unfolds over $1+T$ time steps, starting with an initial task $t_{0}$ and moving through $T$ subsequent tasks, each introducing new sounds for detection. Every task $t_{i}$ has its own dataset $D_{i}$, consisting of sound features $x_{j}^{i}$ and their corresponding labels $y_{j}^{i}$ that are represented as multi-hot vectors. The total number of sound classes $|C|$ evolves to include both existing classes learned in previous tasks ($|C_{\text{ext}}|$) and the new classes introduced in the current task ($|C_{\text{cur}}|$). Our model, designated as $M_{i}$ for each task $t_{i}$, comprises two main components: a feature extractor $\theta_{i}$ and a classifier $\phi_{i}$. The feature extractor analyzes the sound data to identify distinguishing characteristics, while the classifier uses these features to predict the types of sounds present.
II-B Proposed UCIL Method
@@
II-B1 Independent Learning to Update Model by New Data
We start by training a learning model on an initial task, which we refer to as $t_{0}$. This initial task uses a dataset named $D_{0}$. We employ binary cross-entropy loss ($\mathcal{L}_{cls}$) that computes the difference between the model predictions and the ground truth. It operates by applying a function, the sigmoid function $\sigma$, to the model’s logits $o_{k}$ in class $k$ as:
$\mathcal{L}_{cls}=-\frac{1}{|C|}$ $\sum_{k=|C|}^{|C|}y_{k}^{i}\cdot\log(\sigma(o_{k}))$ (1)
@@ II-B1 Independent Learning to Update Model by New Data
After the initial task $t_{0}$, we create an updated version of our model denoted as $M_{i}$ by new sounds in subsequent tasks $t_{i}$. We add the new classification heads of the linear layer, building on the previous model $\hat{M}_{i-1}$ to accommodate the new sound classes $|C_{\text{cur}}|$ of the current task. The predictions made by the model $M_{i}$ are now a mix of existing and new sound classes, which we refer to as $o_{ext}$ and $o_{cur}$, respectively. The updated model, with its new feature extractor $\hat{\theta}_{i}$ and classifier $\hat{\phi}_{i}$, is then trained on the latest data set, $\hat{D}_{i}$. To prevent data imbalance issues of existing and new classes, we use the independent learning (IndL) [22] method, which allows the model to learn the new sound predictions independently from the existing ones. Specifically, we only compute the $\mathcal{L}_{cls}$ by only the new predictions. This way, the existing predictions are maintained separately.
II-B2 Knowledge Distillation from Existing to New
@@
We use two special distillation losses to maintain the learned knowledge of our model [24]. The first one is prediction distillation loss $\mathcal{L}_{dis}^{P}$, which reduces differences in the model’s initial predictions for known sound classes. We use mean squared error (MSE) loss to effectively handle the precise temporal alignment required in strongly labeled SED. The second one is the feature distillation loss $\mathcal{L}_{dis}^{F}$, which minimizes the gap between the current and existing model’s features by comparing their alignment with cosine similarity. The current model $M_{i}$’s feature is the output of the feature extractor $\theta_{i}$ given as $v=\theta_{i}(x)$. In simpler terms, these techniques help our updated model $M_{i}$ act like its existing version, $\hat{M}_{i-1}$, especially for sounds it already learned.
$\mathcal{L}_{dis}^{P}=\lambda MSE(\hat{M}_{i-1}(x)||o_{ext})$ (2)
$\mathcal{L}_{dis}^{F}=1-\cos(\text{norm}(v),\text{norm}(\bar{v}))$
@@
where norm and $\cos$ denote the L2-normalization and cosine similarity, respectively. In detail, we compute the scalar product with normalized vectors to get the cosine similarity. The weight $\lambda=\Omega\sqrt{\frac{|C|}{|C_{cur}|}}$ is set adaptively, $\Omega$ is a fixed constant, and $\lambda$ increases as the number of incremental phases increases, to retain the knowledge of the increasing number of existing classes in the model. For the existing sounds, we measure the output predictions from $M_{i}$ and match them against $\hat{M}_{i-1}$.
@@
Figure 1: Block diagram of the proposed UCIL approach for task $t_{i}$. All training data for task $t_{i}$ include the new data $D_{i}$, the rehearsal data $x^{e}_{i}$, and the unlabeled data $x^{u}_{i}$ as the input. $o_{ext}$, $o_{cur}$, and $o_{u}$ present the prediction of the model $M_{i}$ for the classes of the existing task, classes of the current learning task, and the unlabeled data.
II-B3 Unsupervised Learning with Sample Selection
Given the limited availability of strongly labeled data, researchers have proposed various SED models that incorporate unlabeled data through semi-supervised learning. When unlabeled data influences SED model behavior, it becomes crucial to leverage these predictions to maintain model consistency. Therefore, we introduce unlabeled data $x^{u}_{i}$ into the dataset $D_{i}$ of the task $t_{i}$. However, not all data is helpful—some can confuse the model. Hence, we come up with a novel way to pick the helpful samples from unlabeled data. We look for samples where the model predictions differ greatly due to forgetting, assuming these samples will teach the model effectively. Specifically, we measure the discrepancy between two logit distributions with the L1 metric as:
$d=|\hat{M}_{i-1}(x^{u}_{i})-M_{i}(x^{u}_{i})|\vspace{-2mm}$ (3)
@@ II-B3 Unsupervised Learning with Sample Selection
We consider the first half of samples based on the discrepancy $d$ from our unlabeled data for each batch of training task $t_{i}$. By selecting these samples by discrepancy, we assist the model to retain previously learned knowledge and generalize from the past knowledge to the new situations. We adopt MSE loss given in Section II-B2 to use as $\mathcal{L}_{dis}^{U}$, focusing on unlabeled data. Our goal is to reduce the gap between the new and existing models for the reaction of unlabeled data.
II-B4 Balanced Memory Update Method
@@
To further maintain the previous knowledge, especially in applications with complex classes such as SED, we preserve some of the data from existing classes in advance as the rehearsal data. We mix this data $x^{e}_{i}$ with the incoming data of task $t_{i}$. The balanced sampling function is an algorithm tailored for preparing rehearsal data in the context of SED using strongly and weakly labeled data. It starts by computing the duration of sound events, considering the difference between their onset and offset times. Then, it aggregates these durations for each event, providing a measure of total exposure for each sound in the strongly labeled dataset. For weakly labeled data, the duration is 10 seconds for each event label.
@@
The method makes sure that there is a fair mix of sounds for training the model. It figures out how many examples of each sound are needed based on the seen classes. It then picks a balanced set of sound examples from both types of data to avoid any bias. The outcome is two rehearsal datasets, one from strongly labeled and the other from weakly labeled data. This way, all seen sounds get a fair chance to replay during training, leading to a more reliable and fair SED model.
@@
The approach for independent unsupervised learning as illustrated in Fig. 1. The total loss function $\mathcal{L}_{tot}$ of the updated model is given as:
$\mathcal{L}_{tot}=\mathcal{L}_{cls}+\mathcal{L}_{dis}^{P}+\mathcal{L}_{dis}^{F}+\mathcal{L}_{dis}^{U}\vspace{-2mm}$ (4)
@@
In summary, the model is first trained on one task and then updated with new classifier units for enhanced classification. Independent learning separates the learning of new and existing sound classes. To retain knowledge, we use prediction and feature distillation losses, ensuring consistency across classes. We also strategically select samples from unlabeled data to improve generalization. Finally, our balanced memory update method prepares rehearsal data from both strongly and weakly labeled datasets, ensuring fair representation.
III Experiment Setting
III-A Dataset, Task Setting and Performance Metric
@@
In this study, we use the DCASE 2023 Task 4A dataset [26], a comprehensive collection of 10-second audio clips that are categorized into 10 sound events. It is comprised of 1,578 real recordings with weak annotations, 14,412 real recordings without labels, and 10,000 synthetic recordings strongly labeled [28]. All clips were resampled to 16 kHz mono and segmented using a 2048-sample window and 256-sample hop length for log-mel spectrogram generation.
@@
We devised two incremental task settings to assess our proposed method. Firstly, the two-task setting randomly divides the 10 target sound classes in the dataset into two groups, each containing 5 classes. This approach is designed to test how well our method can adapt from one group of sounds to another, a process known as source-target domain adaptation in SED. Secondly, the four-task setting organizes the 10 target classes into 4 groups randomly where two groups contain two sound events and the rest two contain three sound events. We shuffle the groups as the four tasks to investigate the effect of the number of tasks on the model’s ability to distinguish between different sounds, providing insights into the complexities of learning diverse acoustic events.
@@
III-B Implementation Details and Reference Baselines
We used a batch size of 48 and the Adam [31] optimizer with an initial 0.001 learning rate, gradually increasing over the first 50 of 200 total epochs following the DCASE 2023 Task 4A setup. We adopted a mean teacher model strategy [32] to enhance training stability, applying an exponential moving average with a factor of 0.999. We consider a convolutional recurrent neural network (CRNN) [33] model that serves as the DCASE baseline with about 1.2M parameters for all the studies in this work. We implemented early stopping at 50 epochs to prevent overfitting, and set the weight $\Omega$ for the prediction distillation loss ($\mathcal{L}_{dis}^{P}$) at 2, ensuring a balance between retaining existing knowledge and learning new information. We built following 5 reference baselines:
TABLE I: Performance comparison of our proposed UCIL method with other CIL methods using the two-task settings.